# Forecasting CPI inflation under economic policy and geo-political uncertainties

### v1

@@ 4 Empirical Risk Minimization
FEWNet utilizes wavelet decomposition as a filtering stage and combines it with autoregressive neural networks that incorporate exogenous variables. This combination allows for precise predictions of time series that are both non-stationary and nonlinear. Wavelet decomposition yields a hierarchical structure of new time series derived from the original time series, facilitating their modeling and forecasting. This section demonstrates that FEWNet effectively minimizes the empirical risk in comparison to classical methods, hence highlighting its practical relevance.
@@
$\mathcal{R}_{\text{Gen}}=\lim_{T\to\infty}\frac{1}{T}\sum_{p=1}^{T}\mathbb{E}\left[\left(\hat{Y}_{p+m}-Y_{p+m}\right)^{2}|Y_{p},Y_{p-1},\ldots\right],$
@@
where $\hat{Y}_{i}$ is the prediction (expectations can be omitted for deterministic time series). In practical problems of economic time series forecasting, the relationship between $\hat{Y}_{p+m}$ and the observed time sequence $Y_{p},Y_{p-1},\ldots$ is supposed to be nonlinear in nature. Thus, the inclusion of autoregressive neural networks (a set of feed-forward artificial neural networks designed specifically for predicting future values in time series data) for forecasting is justified in this exercise. The relationship can be estimated by
@@
where $f$ denotes an autoregressive neural network function. The model’s order ($p$) is an essential parameter to estimate. Smaller values of $p$ make the model training simple but miss the information about the past lagged values whereas larger values of $p$ compel the model training hard resulting in a “curse of dimensionality”. To avoid the issues of underfitting and overfitting, we use the ARNNx model, where we choose $p$ by minimizing the SMAPE for the validation set. If $f$ in Eq. (3) is a complex estimator (e.g., deep learning models such as transformers, deep MLP, etc), it will learn well on the training data but perform poorly on unseen test data. Application of MODWT decomposes the time series using low and high pass band filters iteratively resulting in a trend series and several detailed series (containing dynamics of the economic system at different scales). ARNNx is then trained to model the trend and detailed series and provide predictions for future values. The combined values of these MODWT-decomposed series provide a prediction of the original economic data. Although the main motivation for using wavelet decomposition on these economic time series is to enhance the predictive power, MODWT gives the smooth term $S_{\mathcal{K}}$ that contains the slowest dynamics (noise-free) and $\delta_{\tilde{k}}$ that contains system dynamics (noisy). The low-resolution data are frequently contaminated by noise. Consequently, training an ARNNx model on these time series is less challenging compared to the original time series. As $\tilde{k}$ increases, the dynamics get slower, resulting in low-level detail series that consist solely of noise. To mitigate overfitting, one can assign a value of zero to their predictions.
Several recent works [56, 58, 59, 57] have proposed to treat wavelet-decomposed time series independently. Another complicated approach will be to treat each series along with other series (treated as exogenous variables) but it results in increasing the dimensionality of the time series forecasting problem. In the current formulation we model $S_{\mathcal{K}}$ and other $\delta_{\tilde{k}},\;(\tilde{k}=1,2,\ldots,\mathcal{K})$ series using ARNNx models as follows:
@@
The choice of $\hat{f}_{i}(i=0,1,2,\ldots,\mathcal{K})$ is related to the properties of the time series (nonlinear for this case). [57] proposed nonlinear ARNNx model and each $\hat{f}_{i}$ may have the same order for $r_{i}$ (we will use $r$ for the rest) to decrease the model’s complexity. The predictions are made as an equal-weighted combination of smooth and details predictions:
$\hat{Y}_{p+m}=\hat{S}_{\mathcal{K},p+m}+\sum_{\tilde{k}=1}^{\mathcal{K}}\hat{\delta}_{\tilde{k},p+m}.$
@@
Since ARNNx estimators can be expressed as a nonlinear combination of linear projections of lagged inputs for the CPI inflation series, the following result holds:
@@
[…]
Proof 1
Using Eq. (3) we write, the ARNNx estimator as follows
$\hat{f}=\alpha+\sum_{j^{{}^{\prime}}=1}^{q^{*}}\beta_{j^{{}^{\prime}}}\sigma\left(z_{j^{{}^{\prime}}}\right),$
where $z_{j^{{}^{\prime}}}=\alpha^{*}+\sum_{i=1}^{r}\beta_{i,j^{{}^{\prime}}}^{*}y_{i}+\gamma^{*}\underline{X}$ where the parameters are estimated by minimizing Eq. (2). For FEWNet we have
$\begin{gathered}\hat{f}_{0}=\alpha_{0}^{S}+\sum_{i=1}^{q}\beta_{i}^{S}\sigma(\alpha_{i}^{S}+\sum_{j_{0}=1}^{r}{\beta^{S}_{i,j_{0}}}S_{\mathcal{K},j_{0}}+{\gamma^{S}_{i}}^{{}^{\prime}}\underline{X})\\ \hat{f}_{\tilde{k}}=\alpha_{0,\tilde{k}}+\sum_{i=1}^{q}\beta_{i,\tilde{k}}\sigma(\alpha_{i,\tilde{k}}+\sum_{j_{\tilde{k}}=1}^{r}\beta_{i,j_{\tilde{k}}}{\delta}_{\tilde{k},j_{\tilde{k}}}+\gamma^{{}^{\prime}}_{i,\tilde{k}}\underline{X});\;\tilde{k}=1,2,\ldots,\mathcal{K}.\end{gathered}$
[…]
@@ 5 Experimental Evaluation and Analysis of Results
We assessed the efficacy of the proposed FEWNet framework by comparing its performance to other baseline predictions derived from statistical, machine learning, and deep learning methodologies. In our experimental evaluation, we utilize a 5-fold time series cross-validation approach to train all the forecasting frameworks. We then generate forecasts for two different time horizons, specifically 12 months and 24 months, to demonstrate the generalizability of the setup. In this section, we briefly explain the baseline models (Section 5.1), evaluation metrics (Section 5.2), benchmark comparison, and experimental results (Section 5.3).
@@ 5.1 Baseline Models
  • Auto-regressive Integrated Moving Average with exogenous variable (ARIMAx) model is widely used in time series forecasting [60]. The ARIMAx $(p,d,q)$ framework captures the linear patterns of the time series by considering $p$ previous values of the target series, $q$ prior forecast errors, and historical values of the exogenous variable. This framework employs a differencing technique of order $d$ to guarantee the stationarity of the data. The coefficients of the linear ARIMAx model are estimated by minimizing the Akaike information criterion (AIC) value.
@@
  • Deep-learning based Auto-regressive (DeepAR) model is an advanced deep learning system designed specifically for forecasting time series data. This framework employs a recurrent neural network architecture to forecast future trajectories in a dataset that varies over time [61].
@@
  • Auto-regressive Distributed Lag The autoregressive distributed lag (ARDL) model has been extensively employed in the field of econometrics to estimate the long-run and short-run dynamics of time series data [62]. The distributed lag components of the ARDL model are capable of simulating both the stationary and non-stationary time series components.
@@
  • Xtreme Gradient Boosting (XGBoost) is a supervised machine learning algorithm that intends to predict the target variable through the ensemble of a set of weak learners [63]. These tree-based models are trained in a sequential fashion, where the model attempts to minimize the remaining error components of all the prior trees. During prediction, the sum of all the predicted values of the weak learners is reported. In our experimental analysis, we convert the forecasting task into a supervised prediction problem where the lagged inputs and the exogenous variables are the features and the subsequent forecast is the required label for the instance.
@@
  • Wavelet ARIMAx (WARIMAx) is a modified version of the ARIMAx model that use the MODWT decomposition technique for the purpose of predicting time series data [59]. The model breaks down the signal that changes over time into several wavelet and smooth coefficients. It then uses an ARIMAx model to make predictions based on each of these series.The multi-step forecasts are derived by aggregating these candidate forecasts for a desired horizon.
  • Auto-regressive Neural Network with exogenous variable (ARNNx) is a generalization of the classical feed-forward neural network model for the auto-regressive time series process [64]. This model utilizes previous observations of the target time series and the exogenous variables as input to the network. The ARNNx model comprises a single hidden layer positioned between the input and output layers. The ARNNx $(p,k)$ model uses $p$-lagged input values from the input layers and passes them via $k$ neurons in the hidden layer. The value of $k$ is typically calculated by the formula $k=\left[\frac{p+1}{2}\right]$, where $p$ is a given parameter. This approach helps mitigate overfitting and maintain a stable learning architecture [51]. The model gets initialized with random values and trained using the gradient descent back-propagation approach [55].
@@ 5.2 Evaluation Metrics
In our study, we have considered five evaluation metrics to assess the performance of the proposed forecaster and the baseline models in generating the 12-month and 24-month ahead forecasts of the CPI inflation series for the BRIC countries. These metrics include the Root Mean Square Error (RMSE), Mean Absolute Scaled Error (MASE), Symmetric Mean Absolute Percentage Error (SMAPE), Theils’ U1 measure, Median Relative Absolute Error (MDRAE), and Median Absolute Percentage Error (MDAPE). The mathematical expression of these metrics is provided as follows:
@@
[…]
where $y_{t}$ is the ground truth, $\hat{y}_{t}$ is the forecast of the corresponding model, $\hat{y}_{t}^{R}$ is the naive forecast at time $t$, and $m$ is the forecast horizon. By convention, the model with the lowest value of the error metric is considered as the best-performing model [51, 57].
@@ 5.3 Experimental results and baseline comparison
This section focuses on the implementation and effectiveness of the proposed FEWNet architecture in forecasting the Consumer Price Index (CPI) inflation series of the BRIC countries. In order to conduct a thorough assessment, we compare the performance of FEWNet to various cutting-edge algorithms. The FEWNet model is implemented using the R statistical software. The MODWT technique is first implemented using the modwt function from the ‘wavelets’ package. This function decomposes the training data into wavelet and scaling coefficients using the pyramid algorithm with the ‘haar’ filter. The determination of the number of decomposition levels is done conventionally as the floor function of $\log_{e}$(length(train)) [34]. In the subsequent phase, each wavelet (details) and scaling (smooth) coefficient series is modeled with trend and cycle information from the log-transformed EPU, GPRC, and CPI data obtained through economic filters. This is achieved using an ARNNx model fitted with the nnetar function from the ‘forecast’ package in R. To select the number of inputs ($p$) for the individual networks, a time series cross-validation approach is applied with a search range of 1-24. Subsequently, these $p$ inputs are processed through $k=(p+1)/2$ hidden nodes arranged in a single hidden layer. This process generates one-step ahead forecasts for the individual series, which are aggregated to produce the final forecast. The framework iteratively generates multi-step ahead forecasts. The hyperparameters of the FEWNet($p,k$) algorithm used in this study are summarized in Table 3. Furthermore, for implementing classical forecasters such as ARIMAx, SARIMAx, ARDL, ARFIMAx, and WARIMAx, the ‘statsmodels’ library in Python is employed. Deep learning models, including DeepAR, NBeats, and TFT, are implemented using the ‘darts’ library in Python. The machine learning frameworks like Lasso and XGBoost are implemented using the ‘skforecast’ package in Python.
@@
Table 3: Model parameters: $(p,k)$ of FEWNet for BRIC countries across different forecast – horizons. The table shows optimal parameter combinations for 12-month and 24-month rolling window forecasts of CPI inflation.
@@
After implementing the proposed FEWNet model and other baseline architectures, we generate forecasts for the specified time horizons. Tables 4 and 5 display the performance of the models for the semi-long term (12-month) and long-term (24-month) timeframes, respectively, using out-of-sample data. The error metrics reported in the tables indicate that, in the majority of forecasting tasks, the proposed FEWNet framework outperforms the baseline models. In the semi-long term forecasting of Brazil, the FEWNet model significantly reduces the RMSE metric of the ARNNx model. This improvement in model performance is attributed to the application of wavelet decomposition in the FEWNet model. A similar conclusion is supported by other accuracy metrics. For the CPI inflation series of Russia, while the DeepAR model provides the best forecasts among the baseline models, the FEWNet model’s forecasts closely align with the ground truth. The 12-month ahead CPI inflation forecasts of India generated by the proposed architecture show substantial improvement over the best-performing NBeats and WARIMAx models. When forecasting the CPI series of China, the NBeats model competes with the FEWNet architecture, with the former exhibiting the best accuracy for all metrics except SMAPE. The Chinese Consumer Price Index (CPI) series demonstrates a pattern of stable trends (trend-stationary). While the wavelet-decomposition approach is commendable for tackling structural volatility and non-stationarity, its efficacy may be restricted when applied to trend-stationary or covariance stationary series. Consequently, it may not outperform other statistical or machine-learning methods when used on inherently stationary series (a property that is not prevalent for most macroeconomic variables). This explains the cause of the somewhat competitive or less efficient performance of FEWNet in predicting Chinese inflation numbers. In long-term forecasting, the FEWNet architecture consistently outperforms baseline models for all countries except China, where the NBeats model demonstrates more accurate forecasts of the CPI series. Throughout the experimental evaluations, the importance of wavelet transformation and the economic filtered auxiliary information becomes evident in efficiently capturing both time and frequency-level dynamics of the time series. While machine learning and deep learning models generally produce competitive performances for semi-long forecast horizons, their effectiveness diminishes significantly over the long-term horizon. In contrast, the FEWNet model consistently performs well over both horizons, demonstrating its generalizability. Additionally, the wavelet decomposition of the FEWNet approach proves to be a suitable technique for modeling time-frequency information, especially for target variables exhibiting non-stationary behavior (unlike China). Given that most economic time series follow a non-stationary trajectory, the FEWNet approach emerges as an ideal framework for their long-term forecasting even in the presence of structural volatility.
@@
Table 4: Performance of the proposed FEWNet model in comparison to baseline forecasting techniques for 12 months ahead forecasts with exogenous factors EPU, and GPRC (best results are made bold).

### v2

@@ 4 Empirical Risk Minimization
FEWNet utilizes wavelet decomposition as a filtering stage and combines it with autoregressive neural networks that incorporate exogenous variables. This combination allows for precise predictions of time series that are both non-stationary and non-linear. Wavelet decomposition yields a hierarchical structure of new time series derived from the original time series, facilitating their modeling and forecasting. This section demonstrates that FEWNet effectively minimizes the empirical risk compared to classical methods, hence highlighting its practical relevance.
@@
$\mathcal{R}_{\text{Gen}}=\lim_{T\to\infty}\frac{1}{T}\sum_{p=1}^{T}\mathbb{E}\left[\left(\hat{Y}_{p+m}-Y_{p+m}\right)^{2}|Y_{p},Y_{p-1},\ldots\right],$ (5)
@@
where $\hat{Y}_{i}$ is the prediction (expectations can be omitted for deterministic time series). In practical problems of economic time series forecasting, the relationship between $\hat{Y}_{p+m}$ and the observed time sequence $Y_{p},Y_{p-1},\ldots$ is supposed to be non-linear in nature. Thus, the inclusion of autoregressive neural networks (a set of feed-forward artificial neural networks designed specifically for predicting future values in time series data) for forecasting is justified in this exercise. The relationship can be estimated by
@@
where $f$ denotes an autoregressive neural network function. The model’s order ($p$) is an essential parameter to estimate. Smaller values of $p$ make the model training simple but miss the information about the past lagged values, whereas larger values of $p$ compel the model training hard, resulting in a “curse of dimensionality”. To avoid the issues of underfitting and overfitting, we use the ARNNx model, where we choose $p$ by minimizing the SMAPE for the validation set. If $f$ in Eq. (7) is a complex estimator (e.g., deep learning models such as transformers, deep MLP, etc.), it will learn well on the training data but perform poorly on unseen test data. Application of MODWT decomposes the time series using low and high pass band filters iteratively, resulting in a trend series and several detailed series (containing dynamics of the economic system at different scales). ARNNx is then trained to model the trend and detailed series and provide predictions for future values. The combined values of these MODWT-decomposed series provide a prediction of the original economic data. Although the main motivation for using wavelet decomposition on these economic time series is to enhance the predictive power, MODWT gives the smooth term $S_{\mathcal{K}}$ that contains the slowest dynamics (noise-free) and $\delta_{\tilde{k}}$ that contains system dynamics (noisy). The low-resolution data are frequently contaminated by noise. Consequently, training an ARNNx model on these time series is less challenging compared to the original time series. As $\tilde{k}$ increases, the dynamics get slower, resulting in low-level detail series that consist solely of noise. To mitigate overfitting, one can assign a value of zero to their predictions.
Several recent works Aussem & Murtagh (1997); Soltani (2002); Aminghafari & Poggi (2007); Panja et al. (2023) have proposed to treat wavelet-decomposed time series independently. Another complicated approach will be to treat each series along with other series (treated as exogenous variables) but it results in increasing the dimensionality of the time series forecasting problem. In the current formulation we model $S_{\mathcal{K}}$ and other $\delta_{\tilde{k}},\;(\tilde{k}=1,2,\ldots,\mathcal{K})$ series using ARNNx models as follows:
@@
The choice of $\hat{f}_{i}(i=0,1,2,\ldots,\mathcal{K})$ is related to the properties of the time series (non-linear for this case). Panja et al. (2023) proposed non-linear ARNNx model, and each $\hat{f}_{i}$ may have the same order for $r_{i}$ (we will use $r$ for the rest) to decrease the model’s complexity. The predictions are made as an equal-weighted combination of smooth and details predictions:
$\hat{Y}_{p+m}=\hat{S}_{\mathcal{K},p+m}+\sum_{\tilde{k}=1}^{\mathcal{K}}\hat{\delta}_{\tilde{k},p+m}.$ (9)
@@
Since ARNNx estimators can be expressed as a non-linear combination of linear projections of lagged inputs for the CPI inflation series, the following result holds:
@@
Let the autoregressive neural network (of order $r$) is applied to the original data (as in Eq. (7)) by minimizing the risk $\mathcal{R}_{\text{Emp}}$ (as given in Eq. (6)) and FEWNet fits the ensemble model on the decomposed data by minimizing the empirical risk $\mathcal{R}_{\text{Emp}}^{W}$ (as in Eq. (10)), then we have
$\operatorname{min}\mathcal{R}_{\text{Emp}}^{W}\leq\operatorname{min}\mathcal{R}_{\text{Emp}}.$ (11)
Proof. See Appendix 8.4 for the detailed proof.
@@ 5 Experimental Evaluation and Analysis of Results
We assessed the efficacy of the proposed FEWNet framework by comparing its performance to other baseline predictions derived from statistical, machine learning, and deep learning methodologies. In our experimental evaluation, we utilize a 5-fold time series cross-validation approach to train all the forecasting frameworks. We then generate forecasts for two different time horizons, specifically 12 months and 24 months, to demonstrate the generalizability of the setup. In this section, we briefly explain the baseline models (Section 5.1), evaluation metrics (Section 5.2), benchmark comparison and experimental results (Section 5.3), the statistical significance of the results (Section 5.4), empirical validation of theoretical results (Section 5.5), robustness and sensitivity analysis (Section 5.6), and conformal prediction (Section 5.7).
@@ 5.1 Baseline Models
  • Random Walk (RW) is one of the simplest stochastic models based on the assumption that in each period, the time-dependent variable takes a random step away from its previous value, and the steps are independently and identically distributed in size with zero-mean Pearson (1905). The model assumes no mean reversion, which can serve as a naive benchmark. Random Walk with Drift (RWD) is similar to RW but includes a constant drift term.
  • Autoregressive (AR) model forecasts the variable of interest using a linear combination of the past values of the variable. It is like a multiple linear regression with $p$ lagged values of $Y_{t}$ as predictors, thus referred to as AR($p$) model (autoregressive model of order $p$) Hyndman & Athanasopoulos (2018).
  • Markov switching generalized autoregressive conditional heteroscedasticity (MSGARCH) model combines the generalized autoregressive conditional heteroskedasticity (GARCH) model with a Markov-switching process Gray (1996); Haas et al. (2004). This allows for changes in the volatility regime according to a Markov process. MSGARCH accounts for regime changes in volatility, providing a robust benchmark for comparing models in financial econometrics.
  • Autoregressive Integrated Moving Average with exogenous variable (ARIMAx) model is widely used in time series forecasting Box & Pierce (1970). The ARIMAx $(p,d,q)$ framework captures the linear patterns of the time series by considering $p$ previous values of the target series, $q$ prior forecast errors, and historical values of the exogenous variable. This framework employs a differencing technique of order $d$ to guarantee the stationarity of the data. The coefficients of the linear ARIMAx model are estimated by minimizing the Akaike information criterion (AIC) value.
@@
  • Deep-learning based Auto-regressive (DeepAR) model is an advanced deep learning system designed specifically for forecasting time series data. This framework employs a recurrent neural network architecture to forecast future trajectories in a dataset that varies over time Salinas et al. (2020).
@@
  • Auto-regressive Distributed Lag The autoregressive distributed lag (ARDL) model has been extensively employed in the field of econometrics to estimate the long-run and short-run dynamics of time series data Pesaran et al. (1995). The distributed lag components of the ARDL model are capable of simulating both the stationary and non-stationary time series components.
@@
  • Xtreme Gradient Boosting (XGBoost) is a supervised machine learning algorithm that intends to predict the target variable through the ensemble of a set of weak learners Chen & Guestrin (2016). These tree-based models are trained in a sequential fashion, where the model attempts to minimize the remaining error components of all the prior trees. During prediction, the sum of all the predicted values of the weak learners is reported. In our experimental analysis, we convert the forecasting task into a supervised prediction problem where the lagged inputs and the exogenous variables are the features and the subsequent forecast is the required label for the instance.
@@
  • Wavelet ARIMAx (WARIMAx) is a modified version of the ARIMAx model that uses the MODWT decomposition technique for the purpose of predicting time series data Aminghafari & Poggi (2007). The model breaks down the signal that changes over time into several wavelet and smooth coefficients. It then uses an ARIMAx model to make predictions based on each of these series.The multi-step forecasts are derived by aggregating these candidate forecasts for a desired horizon.
  • Auto-regressive Neural Network with exogenous variable (ARNNx) is a generalization of the classical feed-forward neural network model for the auto-regressive time series process Faraway & Chatfield (1998). This model utilizes previous observations of the target time series and the exogenous variables as input to the network. The ARNNx model comprises a single hidden layer positioned between the input and output layers. The ARNNx $(p,k)$ model uses $p$-lagged input values from the input layers and passes them via $k$ neurons in the hidden layer. The value of $k$ is typically calculated by the formula $k=\left[\frac{p+1}{2}\right]$, where $p$ is a given parameter. This approach helps mitigate overfitting and maintain a stable learning architecture Hyndman & Athanasopoulos (2018). The model gets initialized with random values and trained using the gradient descent back-propagation approach Rumelhart et al. (1986).
@@ 5.2 Evaluation Metrics
In our study, we have considered five evaluation metrics to assess the performance of the proposed forecaster and the baseline models in generating the 12 months and 24 months ahead forecasts of the CPI inflation series for the BRIC countries. These metrics include the Root Mean Square Error (RMSE), Mean Absolute Scaled Error (MASE), Symmetric Mean Absolute Percentage Error (SMAPE), Theils’ U1 measure, Median Relative Absolute Error (MDRAE), and Median Absolute Percentage Error (MDAPE). The mathematical expression of these metrics is provided as follows:
@@
where $y_{t}$ is the ground truth, $\hat{y}_{t}$ is the forecast of the corresponding model, $\hat{y}_{t}^{R}$ is the naive forecast at time $t$, and $m$ is the forecast horizon. By convention, the model with the lowest value of the error metric is considered as the best-performing model Hyndman & Athanasopoulos (2018); Panja et al. (2023).
@@ 5.3 Experimental results and baseline comparison
This section focuses on the implementation and effectiveness of the proposed FEWNet architecture in forecasting the CPI inflation series of the BRIC countries. In order to conduct a thorough assessment, we compare the performance of FEWNet to various cutting-edge algorithms. The FEWNet model is implemented using the R statistical software. The MODWT technique is first implemented using the modwt function from the ‘wavelets’ package. This function decomposes the training data into wavelet and scaling coefficients using the pyramid algorithm with the ‘haar’ filter. The determination of the number of decomposition levels is done conventionally as the floor function of $\log_{e}$(length(train)) Percival & Walden (2000). In the subsequent phase, each wavelet (details) and scaling (smooth) coefficient series is modeled with trend and cycle information from the log-transformed EPU, GPRC, and CPI data obtained through economic filters. This is achieved using an ARNNx model fitted with the nnetar function from the ‘forecast’ package in R. To select the number of inputs ($p$) for the individual networks, a time series cross-validation approach is applied with a search range of 1-24. Subsequently, these $p$ inputs are processed through $k=(p+1)/2$ hidden nodes arranged in a single hidden layer. This process generates one-step ahead forecasts for the individual series, which are aggregated to produce the final forecast. The framework iteratively generates multi-step ahead forecasts. The hyperparameters of the FEWNet($p,k$) algorithm used in this study are summarized in Table 4. Furthermore, for implementing classical forecasters such as AR, ARIMAx, SARIMAx, ARDL, ARFIMAx, and WARIMAx, the ‘statsmodels’ library in Python is employed. Deep learning models, including DeepAR, NBeats, and TFT, are implemented using Python’s ‘darts’ library. The machine learning frameworks like Lasso and XGBoost are implemented using the ‘skforecast’ package in Python. RW, RWD have been implemented using ‘forecast’ package in R statistical software and ‘msgarch’ package in R facilitated fitting the MSGARCH model using the FitMCMC function.
Table 4: Model parameters: $(p,k)$ of FEWNet for BRIC countries across different forecast – horizons. The table shows optimal parameter combinations for 12 months and 24 months rolling window forecasts of CPI inflation.
@@
After implementing the proposed FEWNet model and other baseline architectures, we generate forecasts for the specified time horizons. Tables 5 and 6 display the performance of the models for the semi-long-term (12 months) and long-term (24 months) timeframes, respectively, using out-of-sample data. The error metrics reported in the tables indicate that, in the majority of forecasting tasks, the proposed FEWNet framework outperforms the baseline models. In the semi-long-term forecasting of Brazil, the FEWNet model significantly reduces the RMSE metric of the ARNNx model. This improvement in model performance is attributed to the application of wavelet decomposition in the FEWNet model. Other accuracy metrics supports a similar conclusion. For the 12-month forecast horizon, FEWNet consistently outperforms the baseline models for Brazil datasets. In the case of Russia, stochastic AR and FEWNet remained comparative, and their forecasts closely aligned with the ground truth. In the case of India, FEWNet and RWD both outperformed all the benchmark methods by a significant margin. The only stationary time series is the CPI inflation series of China, where the AR model performs the best; however, FEWNet and NBeats remained competitive, and FEWNet obtained the lowest SMAPE amongst its peers. As the Chinese CPI inflation series demonstrates a pattern of steady state (trend-stationary), linear models perform superior to others. While the wavelet-decomposition approach is commendable for tackling structural volatility and non-stationarity, its efficacy may be restricted when applied to trend-stationary or covariance stationary series. Consequently, it may not outperform other statistical or machine-learning methods when used on inherently stationary series (a property that is not prevalent for most macroeconomic variables). This explains the cause of the somewhat competitive or less efficient performance of FEWNet in predicting Chinese inflation numbers. In long-term forecasting, the FEWNet architecture consistently outperforms baseline models for all countries except China and Russia. In the case of China, NBeats, DeepAR, and Lasso regression demonstrate more accurate forecasts of the CPI inflation series. For the Russia CPI inflation series, the AR model outperformed FEWNet and other benchmark methods. Throughout the experimental evaluations, the importance of wavelet transformation and the economically filtered auxiliary information becomes evident in efficiently capturing both time and frequency-level dynamics of the time series. While machine learning and deep learning models generally produce competitive performances for semi-long forecast horizons, their effectiveness diminishes significantly over the long-term horizon. In contrast, the FEWNet model consistently performs well over both horizons, demonstrating its generalizability. Additionally, the wavelet decomposition of the FEWNet approach proves to be a suitable technique for modeling time-frequency information, especially for target variables exhibiting non-stationary behavior (unlike China). Given that most economic time series follow a non-stationary trajectory, the FEWNet approach emerges as an ideal framework for their long-term forecasting even in the presence of structural volatility.
@@
Table 5: Performance of the proposed FEWNet model in comparison to baseline forecasting techniques for 12 months ahead forecasts with exogenous factors EPU and GPRC (best results are made bold).